Two judges in a beauty contest rank the 12 entries as follows

12 Y 12 9 6 10 3 5 4 7 8 2 11 1

Calculate the rank correlation coefficient between X and Y.


1
Expert's answer
2021-02-22T13:21:52-0500

The enteries by two judges are:-


Here "N=12" ,

Rank correlation coefficient ="1-\dfrac{6\times \sum D^2}{N^3-N}"


="1-\dfrac{6\times 416}{(12)^3-12}"


="1-\dfrac{6\times 416}{1728-12}"


=1-"\dfrac{6\times 416}{1716}"

="1-\dfrac{416}{280}=1-1.45=-0.45"


Hence Rank correlation coefficient is "-0.45."